OpenStudy (anonymous):

Find the center of the circle defined by the equation. x^2+y^2+16x+10y+53=0 standard form of the equation of a circle is: (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Complete the square.

OpenStudy (anonymous):

complete the square to write it in that form

OpenStudy (anonymous):

do you know how to complete the square?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

\[x^2+16x+y^2+10y=-53\]

OpenStudy (anonymous):

i just grouped the terms together. now half of 16 is 8 and half of 10 is 5 so write \[(x+8)^2+(y+5)^2=-53+8^2+5^2\]

OpenStudy (anonymous):

the reason this works is that \[(x+8)^2=x^2+8x+16\] so you have added 16 to the left and therefore have to add 16 to the right

OpenStudy (anonymous):

check this: (x+8)^2 -64+ (y+5)^2 -25 +53=0 (x+8)^2 +(y+5)^2=36 here is your center: (-8,-5)

OpenStudy (anonymous):

likewise \[(x+5)^2=x^2+10x+25\] so you are adding 25 when you rewrite in that fashion. adjust by adding 25 to the right

OpenStudy (anonymous):

Is it (3,5)

OpenStudy (anonymous):

no center is (-3,-5)

OpenStudy (anonymous):

not that either

OpenStudy (anonymous):

(-8,-5) ?

OpenStudy (anonymous):

you get \[(x+8)^2+(y+5)^2=36\] yes?

OpenStudy (anonymous):

so center is (-8,-5) and radius is 6

OpenStudy (anonymous):

that's that i put up there... :)

OpenStudy (anonymous):

x^2 + y^2 - 6x - 10y + 30 = 0 .. x^2-6x +y^2-10y+30=0 (x^2-6x+9)-9+(y^2-10y+25)-25+30=0 (x-3)^2+(y-5)^2-9-25+30=0 (x-3)^2+(y-5)^2=9+25-30 (x-3)^2+(y-5)^2=4 Center = (3,5) sum = 8

OpenStudy (anonymous):

this may be correct but it is not the problem up top!

OpenStudy (anonymous):

agreed with satellite73 :)

OpenStudy (anonymous):

I I messed the 16 up! uggghh.

OpenStudy (anonymous):

btw i find it easier to just put all the numbers on the right hand side instead of adding and subtracting. of course it doesn't matter, it is just my personal choice.

OpenStudy (anonymous):

in other words if i see \[x^2+10x-3=0\] i would rewrite as \[x^2+10x=3\] and then put \[(x+5)^2=3+25